OpenStudy (anonymous):

What is the slope of the line that passes through 4, 1 and 7, -2? 1/3 1 -1 -1/3?

pooja195 (pooja195):

Hi welcome to opentudy :) Use the slope formula \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

pooja195 (pooja195):

Are you familiar with that?

OpenStudy (anonymous):

Yes but idr how to do it

pooja195 (pooja195):

Ok can you start by tell me what: y1=? y2=? x1=? x2=? Can you tell which values correspond with each?

pooja195 (pooja195):

(4, 1) ( 7, -2) (x1,y1)(x2,y2) See if this helps

OpenStudy (anonymous):

Okay, I get 5?

pooja195 (pooja195):

How did you get that?

OpenStudy (anonymous):

4-1/7- -2

pooja195 (pooja195):

Not quite the way ordered pairs work are like this (x,y) (x1,y1) So we have: x1=4 x2=7 y1=1 y2=-2 \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\] \[\LARGE \frac{-2-1}{7-4}=slope\] Now solve and find the slope :)

OpenStudy (anonymous):

Ohh ok, -1?

pooja195 (pooja195):

yes well done :)

OpenStudy (anonymous):

Thank you! :)

pooja195 (pooja195):

yw ^_^ \(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)